OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6.

OpenStudy (anonymous):

when you have ax + by = c you find the x-int by c/a and the y-int c/b In you problem a = 3, b = -2, c = 6 so x -int = 6/3 = 2 and y-int = 6/-2 = -3